OpenStudy (anonymous):

Help? Kerry is planning to plant a rectangular garden that is 16 feet longer than it wide. Kerry wants to find its width if its area is 225 square feet. Which equation models this problem? (Points : 1) w + 16w = 225 w2 – 16w = 225 w2 + 16w = 225 16w2 = 225

OpenStudy (anonymous):

A. w + 16w = 225 B. w2 -16 w = 225 C.w2 + 16w =225 D 16w2 = 225